Canadian Bacon Inc. financial statements are presented in the table below.

Based on the information in the table, and using a 365-day year, calculate cash conversion cycle

Round the answers to two decimal places

Balance Sheet December 31, 2012

Cash and marketable securities $198,000 Accounts payable $288,000
Accounts receivable $469,000 Notes payable $65,000
Inventories $577,000 Accrued expenses $84,000
Prepaid expenses $15,700 Total current liabilities $437,000
Total current assets $1,259,700 Long-term debt $237,000
Gross fixed assets $1,954,000 Par value and paid-in-capital $199,000
Less: accumulated depreciation $476,000 Retained Earnings $1,864,700
Net fixed assets $1,478,000 Common Equity 2,063,700
Total assets $2,737,700 Total liabilities and owner’s equity $2,737,700

Income Statement, Year of 2012

Net sales (all credit) $7,546,600.00
Less: Cost of goods sold $6,112,746.00
Selling and administrative expenses $349,000.00
Depreciation expense $145,000.00
EBIT $939,854.00
Interest expense $49,500.00
Earnings before taxes $890,354.00
Income taxes $356,141.60
Net income $534,212.40
Answer #1

purchases = CoGS-payables = 6112746-288000 =

5824746

Cash cycle = 365 * avg receivables/net credit sales + 365 * avg inventory/COGS- 365*avg payables/purchases

=365*469000/7546600+365*577000/6112746-365*288000/5824746 = 39.09 days